OpenStudy (anonymous):

the proportion is blank over 4 = 21 over 12 and i have to solve it.

OpenStudy (anonymous):

so you have blank/4=21/12 what do you have to take 4 times to get 12, answer 3, or 12/3 = 4 so 21/3 = 7 the blank is 7

OpenStudy (anonymous):

wow thank you

OpenStudy (anonymous):

6 over 10 equals 9 over y i need to solve it

OpenStudy (anonymous):

wer're going to do this one a little different cross multiply: 6 * y = 10 * 9 6*y = 90 divide both sides by 6, so 90/6 = 15

OpenStudy (anonymous):

come again?

OpenStudy (anonymous):

im not good at math

OpenStudy (anonymous):

you have 6 over 10 = 9 over something or 6/10 = 9/box take 9/6 = 3/2 reduced then take 10 *(3/2) = (10*3)/2=30/2=15

OpenStudy (anonymous):

so the blank is 15?

OpenStudy (anonymous):

you're right

OpenStudy (anonymous):

oh okay